The product of two consecutive odd numbers is 143. Find the numbers. (Hint: if the first odd number is x, what is the next odd number?)
Found 2 solutions by swincher4391, solver91311:
Answer by swincher4391(1107)

Let x be the first odd number.
Then the second odd number is x+2 since there is a common difference of 2 between all odd numbers.
Then x(x+2) = 143
x^2 + 2x = 143
x^2 + 2x -143 = 0
What multiplies to -143 and adds to 2?
It may take a while to think of the answer, but let's just think about this for a second.
We need two numbers that add up to 2.
Start with -10 and 12 which is -120... too low
-11 and 13? -143... perfect.
So we get (x-11) and (x+13) Which gives us x =11 and x=-13. It's not likely that our answer is negative, so let's stick with x =11.
Then our second odd number is 11+2 = 13
So our numbers are 11 and 13.

Answer by solver91311(24713)



The next consecutive odd integer is , so





Solve for and discard the negative root. Then calculate

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism